OpenStudy (anonymous):

simplify: radical 27 + radical 75

OpenStudy (hoblos):

radical 27 = 3radical3 radical 75 = 5radical3 radical 27 + radical 75 = 8 radical 3

OpenStudy (anonymous):

Recall \[\sqrt{ab}=\sqrt{a}\sqrt{b}\], so \[\sqrt{27}+\sqrt{75}=\sqrt{9}\sqrt{3}+\sqrt{25}\sqrt{3}=3\sqrt{3}+5\sqrt{3}=8\sqrt{3}\]